OpenStudy (anonymous):

Write an expression to describe a rule for the sequence. Then find the 100th term in the sequence. 4, 9, 14, 19, 24, 29, . . .

OpenStudy (ja1):

+5

OpenStudy (ja1):

so now you basically add 5 all the way to the 100th number

OpenStudy (anonymous):

A. n + 5; 504 B. 5n − 1; 499 C. 5n; 500 D. 4n + 5; 405

OpenStudy (ja1):

ahh much better

OpenStudy (hunus):

Try subtracting the nth term from the (n+1)th term to see if it's incrementing by the same amount

OpenStudy (ja1):

well the only one that says +5 is A. but do you know why the answer is A.?

OpenStudy (anonymous):

still dnt get it

OpenStudy (ja1):

well lets take a look here: 4, 9, 14, 19, 24, 29 How do you get from 4 to 9? you add 5 right, and same with all the rest

OpenStudy (anonymous):

so u add by 5 okay i get it now thank you

OpenStudy (ja1):

kk :)

OpenStudy (anonymous):

D = 2zv for v

OpenStudy (anonymous):

A. v equals 2z over D B. v equald D over 2z C. v = D – 2z D. v equals D over z

OpenStudy (anonymous):

Did you find the general term yet??

OpenStudy (anonymous):

Do you think A is right answer in previous question you asked?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Certainly not..

OpenStudy (anonymous):

Here, you can see every term has a difference of 5, First term you are given with is a = 4 and d, then common difference is 5.. So, this is forming an Arithmetic Progression, general term of which is given by: \[a_n = a+ (n-1)d\] Here : \(a = 4\) and \(d = 5\).. Can you plug in the values and find \(a_n\) here?

OpenStudy (anonymous):

i dnt get it

OpenStudy (anonymous):

See, here you are asked to find the general term first and using that you are to find 100th term.. Now, in this formula: \[a_n = a + (n-1)d\] \(a_n\) is your general term or you can say the nth term. \(a\) is your first term in the sequence, (here it is 4) \(n\) is number of terms. \(d\) is common difference. (In this case : 5) So, these values you have to substitute in that formula, you will get: \[a_n = 4 + (n-1)(5) \implies a_n = 4 + 5n - 5 \implies \color{green}{a_n = 5n-1}\]

OpenStudy (anonymous):

So, this is your general term.. Now can you find 100th term here? Hint : you have to just put n = 100 in the following general term you have got: \[a_n = 5n - 1\] Here just put \(n = 100\) and find \(a_n\).

OpenStudy (anonymous):

General term means \(a_n\) that in turn means "Rule for a Particular Sequence". Here : \(a_n = 5n - 1\) is Rule of the sequence you are given with..

OpenStudy (anonymous):

Getting @Dyalis_love ?? See, if you have doubt or problem anywhere, just ask me and I will help whatever extent I can be..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

If you are getting then tell me what is 100th term here, just put n = 100 in the general term you have got..

OpenStudy (anonymous):

thx for helping me i have to go but thx =)

OpenStudy (anonymous):

Sure..

OpenStudy (anonymous):

I can see you have closed the question but it is fine with me.. Searching for only answers is no wisdom here, you have to understand how we solve each and every problem in Mathematics..